# Unveiling the Flaws: A Critical Analysis of Initialization Effect on Time Series Anomaly Detection

### v1

@@ Abstract
Time-Series Anomaly Detection Anomaly Detection Temporal Learning Reliable Machine Learning.
1 Introduction
Time series anomaly detection (TSAD) has been a significant area of interest in the field of data science due to its significant applications across various domains, including transportation [12], healthcare [8], and industrial monitoring [9]. The recent surge in interest is partly driven by the advancements in deep learning, which have shown promising results in handling time series data for tasks such as classification and forecasting [21].
In recent years, most research in time-series anomaly detection focused on modifying visual anomaly detection frameworks by incorporating a customized temporal architecture to handle sequential data. Early experiments have hinted that these algorithms can significantly outperform the state-of-the-art [4]. However, recent studies have cast a shadow of doubt over these claims [6]. They have shown that the performance of deep time-series models is not necessarily better than traditional time-series anomaly detection algorithms or even, in some cases, random baselines [13]. The superior performance that some papers have reported can be attributed to other factors, such as incorrect evaluation procedures and inappropriate benchmark detests [22]. As a result, most of the recent progress in deep time-series anomaly detection is practically useless in many real-life applications.
Initialization is one of the key aspects of time-series analysis that potentially plays an important role in the performance and reliability of TSAD algorithms. It can lead to misleading conclusions about the efficacy of TSAD algorithms, as the initialization steps can significantly affect the anomaly detection results. This phenomenon challenges the reliability of a TSAD method, as it makes it difficult to distinguish between genuine performance improvements due to the model and variations in results induced by initialization.
Despite the importance of initialization, its impact on anomaly detection performance has not been thoroughly investigated. In this paper, we provide a critical analysis of how initialization affects TSAD performance. We aim to uncover the flaws and biases introduced by commonly employed initialization and preprocessing techniques. Through extensive experiments, we demonstrate how different initialization techniques can significantly alter detection outcomes, highlighting the risk of artificial performance boosts resulting from variations in initialization.
Our findings suggest that while initialization is an integral part of time series processing, it must be applied judiciously, with a clear understanding of its potential impact on anomaly detection results. By shedding light on the complexities and pitfalls of initialization in time series anomaly detection, we hope to contribute to the development of more robust and reliable detection methodologies, ultimately advancing the field and its practical applications.
2 Background
2.1 Anomaly Types
The primary goal of TSAD is to detect anomalies that deviate from the expected data patterns, which can be classified into three main types [11]:
2.1.1 Contextual Anomalies:
These anomalies occur when data points significantly differ from the normal pattern based on the context, such as seasonal trends or temporal dependencies. For instance, unusually high electricity usage during a typically low-demand period could be a contextual anomaly.
2.1.2 Collective Anomalies:
These involve a series of data points that, when considered together, deviate from the normal pattern. An example could be a sudden spike in network traffic that, while individual data points may not seem anomalous, collectively indicate a potential security breach.
2.1.3 Point Anomalies:
These are individual data points that significantly deviate from the expected value. Point anomalies are very common in current TSAD datasets and can include instances like a sudden drop in stock prices.
2.2 Time-Series Anomaly Detection Methods
TSAD has been a longstanding research problem. Several types of methods have been developed to tackle it in the past literature:
2.2.1 Statistical Methods:
These include techniques like Z-score, moving averages, and Holt-Winters exponential smoothing [18]. These methods are straightforward and computationally inexpensive but often lack robustness against complex and high-dimensional data. They assume that data follows a specific statistical distribution, which is not always the case in real-world scenarios.
2.2.2 Traditional Machine Learning Methods:
Unsupervised methods such as Isolation Forest [16], One-Class SVM and SVDD [10], and clustering-based approaches [19]. These methods learn the normal behaviour from training data and identify deviations in test data. They do not require labelled data, making them suitable for real-world applications where anomalies are rare and difficult to label.
2.2.3 Deep Learning Methods:
Techniques like Autoencoders, LSTM networks, and Generative Adversarial Networks (GANs) have shown great promise in TSAD [4]. These methods can model complex temporal dependencies and detect anomalies based on reconstruction errors (e.g., Autoencoders) or prediction deviations (e.g., LSTMs). They require significant computational resources and are often data-hungry, but their ability to capture non-linear relationships in the data makes them powerful tools for TSAD.
2.3 Challenges in Time Series Anomaly Detection
Despite the recent advancements in TSAD, several issues remain unresolved, leading to challenges in accurately evaluating and comparing different anomaly detection methods:
2.3.1 Evaluation Metrics:
Traditional point-based metrics (e.g., precision, recall, F1-score) often fall short of accurately reflecting the performance of TSAD methods, especially for range-based anomalies. Most anomalies in time series data appear consecutively, forming anomaly segments. Point-based evaluation metrics evaluate each observation independently, which can lead to considerable underestimation or overestimation of detection performance.
Recent studies have proposed adjusted evaluation protocols to address these limitations. Point Adjustment (PA) [17] is one such protocol where all points within an anomaly segment are considered true positives if any point within the segment is detected. However, PA can significantly overestimate performance and create an illusion of superior detection capability. Variants like PA%K [13] apply PA only if the ratio of correctly detected anomalies to the segment length exceeds a threshold, aiming to balance precision and recall more effectively. Despite these efforts, the search for a universally effective evaluation metric continues.
2.3.2 Benchmark Datasets:
Commonly used benchmark datasets for TSAD, such as Yahoo [15] and Numenta [2], have been found to have several flaws, including being too simplistic or not representing the complexities of real-world scenarios [22]. These datasets often contain artificially generated anomalies, which do not capture the subtlety and diversity of anomalies encountered in practice. Consequently, methods that perform well on these benchmarks may fail when applied to real-world data, leading to an illusion of progress in the field of TSAD. There is a need for more comprehensive and challenging datasets that better represent real-world conditions to truly evaluate and advance TSAD methods.
2.3.3 Model Interpretability:
Deep learning models, while powerful, often operate as black boxes, making it challenging to interpret their predictions. Understanding why a model flagged certain data points as anomalies is crucial for trust and adoption in critical applications. Methods like LIME and SHAP [1] have been developed to provide explanations for model predictions, but these techniques are still in their infancy for time series data and need further refinement to be effectively applied in TSAD.
2.3.4 Initialization Effect:
Initialization steps such as preprocessing, seed variability, and window size selection are crucial for Time Series Anomaly Detection (TSAD). However, these steps can potentially introduce biases that affect anomaly detection performance. Some initialization steps might create shortcuts for the model, compromising the robustness of the detection process. Therefore, understanding the potential effects of each preprocessing step is essential to ensure the reliability of TSAD methods. Without this understanding, a method might falsely claim to improve state-of-the-art (SOTA) results by inflating its performance through choosing a specific seed or hyperparameters.
By unveiling the flaws and biases introduced by common initialization techniques, we seek to provide a more nuanced understanding of their effects on the accuracy and reliability of anomaly detection methods. Through extensive experiments on benchmark datasets, we aim to demonstrate how different initialization steps can significantly alter detection outcomes. In summary, our contributions are as follows:
  1. 1.
    We evaluate various initialization techniques and their impact on the performance of TSAD methods.
  2. 2.
    We demonstrate how significantly the choice of hyperparameters can impact performance. In many cases, the variation of results caused by altering a benign hyperparameter such as the seed number is far greater than the claimed improvements over the SOTA.
  3. 3.
    By shedding light on this overlooked area of TSAD, we aim to encourage researchers to critically analyze the results of previous work and conduct experiments in future method developments to ensure that performance improvements are due to the model, not the choice of initialization parameters.
3 Experiment Setup
@@
In order to demonstrate the effect of various initialization techniques on the performance of TSAD methods, we designed a comprehensive experimental setup. This involved systematically varying key initialization parameters on benchmark datasets.
3.1 Datasets
3.1.1 Secure Water Treatment (SWaT) Dataset [7]:
@@
3.1.2 Server Machine Dataset (SMD) [20]:
@@
3.2 Methods
For our experiment, we used three state-of-the-art (SOTA) methods that have been commonly employed by TSAD researchers for comparison and benchmarking over the past few years.
3.2.1 Graph Deviation Network (GDN):
GDN [5] is designed to detect anomalies in multivariate time series data by leveraging the inherent structure of the data. The idea behind GDN is to model the relationships between different time series using a graph structure where each node represents a time series, and edges capture the dependencies between them. GDN uses Graph Neural Networks (GNNs) to learn these dependencies and identify deviations from normal behavior. In GDN, each node corresponds to a sensor, and the edges are weighted based on the correlation between the sensors. The network is trained to predict the value of each sensor at a given time step based on the values of the other sensors and itself in the past. Anomalies are detected when the actual sensor values significantly deviate from the predicted values. This method allows GDN to capture and leverage the spatial dependencies in multivariate time series data.
3.2.2 Multivariate Time-Series Anomaly Detection via Graph Attention Networks (MTAD-GAT):
MTAD-GAT [23] employs a combination of convolutional layers, graph attention networks, and recurrent units to capture both spatial and temporal dependencies in multivariate time series data. The architecture begins with a 1-D convolutional layer that extracts high-level features from each time series, enabling the network to understand the inherent patterns within the data. Following this, two parallel Graph Attention Networks (GAT) layers are utilized: one focusing on feature-oriented dependencies and the other on time-oriented dependencies. This dual approach enables the model to capture the intricate relationships between different features and across different time steps. The outputs from the convolutional and GAT layers are then fed into a Gated Recurrent Unit (GRU), which models the sequential patterns in the data, providing a deeper understanding of temporal dependencies. MTAD-GAT uses both a forecasting model, which predicts the next time step, and a reconstruction model, which employs a Variational Autoencoder (VAE) to learn a latent representation of the time series.
3.2.3 Unsupervised Anomaly Detection (USAD):
@@
4 Results and Discussion
4.1 Window Size Effect
@@
A small window size captures fine-grained temporal patterns but may miss broader trends and context. Conversely, a larger window size captures broader temporal trends but may smooth over important short-term anomalies. The choice of window size should balance the need to capture relevant patterns while avoiding the dilution of critical anomalies. Ideally, a reliable method should not be hypersensitive to this parameter. While the previous discussion is valid for very large or very small window sizes, minor adjustments in window size should not significantly impact performance in an ideal scenario. To assess the effect of window size on the anomaly detection performance of our TSAD methods, we set it to 1, 3, 5, 10, 20, 50, 100, and 200 to represent a wide range of possible values. The resulting figure is plotted in Figure 1.
@@
[…]
Not only is there a lack of correlation between window size and performance, but using a window size of 1 actually yielded the highest F1 score in this experiment. This is unexpected because GDN is a predictive method, and a window size of 1 means the model only analyzes the previous time point when determining if a certain time point is anomalous. This suggests that GDN on SWAT is not effectively capturing long-term dependencies when using larger window sizes; if it were, we would expect better performance for window sizes above 1, at least up to a certain point. We performed similar window size experiments on the MTAD-GAT and USAD models and observed the same trend, as discussed in the Supplementary Material section.
@@
Our main takeaway is that these models do not sufficiently utilize their window size architecture to capture long-term dependencies in datasets such as SWAT. Additionally, they suffer from large performance variance relative to the window size, often much more than 5%. This variation in performance is far greater than what current methods achieve over traditional methods, challenging the reliability of their claims, as the observed performance boost might be attributed to changing a parameter such as window size. It is also worth noting that there is no consistency in the literature regarding the standard window size for a dataset. Furthermore, some papers claim to downsample the data due to memory concerns, which also impacts the amount of information in each window and can drastically affect performance compared to the existing SOTA methods.
4.2 Seed Variability
@@
Many model initialization and data preprocessing steps involve randomness. To ensure the reproducibility and reliability of results, it is crucial to evaluate the impact of seed variability. Running experiments with different seeds helps in understanding the stability of the initialization methods and the algorithms’ performance. Ideally, changing the seed number when running a method should not significantly affect performance. If it does, it suggests that the method is not reliable and that optimal seed selection might have been done using grid search to inflate the model’s performance compared to established baselines.
In our experiments, we ran the models with their default parameters but fixed the seed to various seed numbers. We chose seed numbers $\{1,2,\ldots,25\}{ 1 , 2 , … , 25 }$ in these experiments, although any $2525$ different numbers would work. The goal was to examine the variability of the models.
Figure 2: F1 scores obtained by GDN on the SWAT dataset for varying random seed numbers.
@@
Figure 3: F1 scores obtained by MTAD-GAT on the SWAT dataset for varying random seed numbers.
[…]
Figure 3 displays the same trend, this time with the MTAD-GAT model: [^1] In these experiments the anomaly threshold is set to the maximum anomaly score in the validation set, similar to the other experiments and to GDN’s implementation. Some trials yield F1 scores above $0.80.8$ while others yield F1 scores below $0.30.3$. We observe once again that with no input parameter change, there is a discrepancy between the F1 and overall performance of MTAD-GAT across different trials.
@@
The results of these experiments reiterate the point raised in the previous section: current SOTA models seem to be overly sensitive to hyperparameters such as seed number or window size. This sensitivity can provide a shortcut to improved performance simply by searching for and reporting the best parameters for the model rather than for the baselines. We observe performance changes as high as double digits just by perturbing the window size or changing the seed number, which is far greater than the claimed improvements due to their innovative architecture compared to SOTA. This casts a shadow of doubt over which factor plays the major role in the observed improvements: the model itself or the initialization parameters.
4.3 Normalization
@@
4.4 Analysis of Anomaly Score
@@
As observed in previous sections, many seemingly unrelated parameters can significantly affect the performance of a TSAD model. As several studies have hinted earlier, the current numerical improvements reported by many papers are often owed to flawed thresholding and evaluation protocols . To have a more realistic view of the true output of TSAD models under different conditions, we analyze the density of raw anomaly scores before applying any transformation, thresholding, or point adjustment. We plot the density for training, validation, and test sets: in the SWAT and SMD datasets, the training and validation sets consist only of normal data, while the test set includes both normal and abnormal data. Ideally, we would like to observe significantly low anomaly scores for the training, validation, and normal test data, and significantly higher anomaly scores for the abnormal test data.
Figure 5 illustrates the anomaly score density of GDN on the SMD dataset. Densities of other methods and datasets are also plotted in the supplementary materials. In this figure, we can observe that the anomaly scores of the training samples are what we expect: near zero, since the model learned during training that these were normal. The validation set, which consists of unseen normal data, also shows anomaly scores close to zero.
However, the density of the anomaly scores for abnormal samples during the test phase does not significantly differ from the normal data, and is still close to zero. This explains why the model does not perform well in this experimental setup: it cannot distinguish between normal and abnormal samples. Yet, by incorporating schemes such as point adjustment, it might be possible to report an acceptable F1 score.
Another interesting and extreme example of counterintuitive anomaly score densities is observed with MTAD-GAT on the SWAT dataset. As Figure 6 demonstrates, the anomaly scores of the normal test data are even larger than those of the abnormal data. This is counterintuitive given that the model is trained to minimize the anomaly score on normal data. However, since the density of normal test data scores does not fully overlap with the distribution of abnormal anamaly scores, the model could still achieve a good F1 score by using a kernel to transform the anomaly scores and thresholding to isolate the densities. This highlights how post-processing techniques can mask underlying model deficiencies and produce seemingly favorable results.
At the cost of sacrificing the robustness of the model, one could thoroughly search for a set of optimal initialization parameters such as window size, seed number, and normalization technique. Combined with existing flawed evaluation protocols, this approach could potentially outperform any SOTA method without actually improving the model’s impact on real-world TSAD problems. This underscores the need for rigorous evaluation and transparent reporting of preprocessing steps and input data characteristics to ensure the reliability and fairness of anomaly detection methods.
Figure 5: Density plot of anomaly scores obtained by GDN on the SMD dataset. The plot illustrates the distribution of anomaly scores across different sets: the training set, validation set, normal data from the test set, and abnormal data from the test set.
@@
Figure 6: Density plot of anomaly scores obtained by MTAD-GAT on the SWAT dataset. The plot illustrates the distribution of anomaly scores across different sets: the training set, validation set, normal data from the test set, and abnormal data from the test set.
5 Conclusion

### v2

@@ Abstract
Time-Series Anomaly Detection, Anomaly Detection, Temporal Learning, Reliable Machine Learning.
I Introduction
Time series anomaly detection (TSAD) has been a significant area of interest in the field of data science due to its significant applications across various domains, including transportation [1], healthcare [2], and industrial monitoring [3]. The recent surge in interest is partly driven by the advancements in deep learning, which have shown promising results in handling time series data for tasks such as classification and forecasting [4]. In recent years, most research in time-series anomaly detection focused on modifying visual anomaly detection frameworks by incorporating a customized temporal architecture to handle sequential data. Early experiments have hinted that these algorithms can significantly outperform the state-of-the-art [5]. However, recent studies have cast a shadow of doubt over these claims [6]. They have shown that the performance of deep time-series models is not necessarily better than traditional time-series anomaly detection algorithms or even, in some cases, random baselines [7]. The superior performance that some papers have reported can be attributed to other factors, such as incorrect evaluation procedures and inappropriate benchmark detests [8]. As a result, most of the recent progress in deep time-series anomaly detection is practically useless in many real-life applications. Initialization is one of the key aspects of time-series analysis that potentially plays an important role in the performance and reliability of TSAD algorithms. It can lead to misleading conclusions about the efficacy of TSAD algorithms, as the initialization steps can significantly affect the anomaly detection results. This phenomenon challenges the reliability of a TSAD method, as it makes it difficult to distinguish between genuine performance improvements due to the model and variations in results induced by initialization. Despite the importance of initialization, its impact on anomaly detection performance has not been thoroughly investigated. In this paper, we provide a critical analysis of how initialization affects TSAD performance. We aim to uncover the flaws and biases introduced by commonly employed initialization and preprocessing techniques. Through extensive experiments, we demonstrate how different initialization techniques can significantly alter detection outcomes, highlighting the risk of artificial performance boosts resulting from variations in initialization. Our findings suggest that while initialization is an integral part of time series processing, it must be applied judiciously, with a clear understanding of its potential impact on anomaly detection results. By shedding light on the complexities and pitfalls of initialization in time series anomaly detection, we hope to contribute to the development of more robust and reliable detection methodologies, ultimately advancing the field and its practical applications.
II Experiment Setup
@@
To demonstrate the effect of various initialization techniques on the performance of TSAD methods, we designed a comprehensive experimental setup which involved systematically varying key initialization parameters on benchmark datasets.
II-A Datasets
II-A1 Secure Water Treatment (SWaT) Dataset [9]
@@
II-A2 Server Machine Dataset (SMD) [10]
@@
II-B Methods
For our experiment, we used three state-of-the-art (SOTA) methods commonly employed by TSAD researchers for comparison and benchmarking over the past few years.
II-B1 Graph Deviation Network (GDN)
GDN [12] is designed to detect anomalies in multivariate time series data by leveraging the inherent structure of the data. The idea behind GDN is to model the relationships between different time series (sensors) using a graph structure where each node represents a time series, and edges capture the dependencies between them. GDN uses Graph Neural Networks (GNNs) to learn these dependencies and identify deviations from normal behavior. The network is trained to predict the value of each sensor at a given time step based on the values of the other sensors and itself in the past. Anomalies are detected when the actual sensor values significantly deviate from the predicted values. This method allows GDN to capture and leverage the spatial dependencies in multivariate time series data.
II-B2 Multivariate Time-Series Anomaly Detection via Graph Attention Networks (MTAD-GAT)
MTAD-GAT [13] employs a combination of convolutional layers, graph attention networks, and recurrent units to capture both spatial and temporal dependencies in multivariate time series data. The architecture begins with a 1-D convolutional layer that extracts high-level features from each time series, enabling the network to understand the inherent patterns within the data. Following this, two parallel Graph Attention Networks (GAT) layers are utilized: one focusing on feature-oriented dependencies and the other on time-oriented dependencies. This dual approach enables the model to capture the intricate relationships between different features and across different time steps. The outputs from the convolutional and GAT layers are then fed into a Gated Recurrent Unit (GRU), which models the sequential patterns in the data, providing a deeper understanding of temporal dependencies.
II-B3 Unsupervised Anomaly Detection (USAD)
@@
III Results and Discussion
III-A Window Size Effect
@@
A small window size captures fine-grained temporal patterns but may miss broader trends and context. Conversely, a larger window size captures broader temporal trends but may smooth over important short-term anomalies. The choice of window size should balance the need to capture relevant patterns while avoiding the dilution of critical anomalies. Ideally, a reliable method should not be hypersensitive to this parameter: very minor adjustments in window size should not significantly impact performance in an ideal scenario. To assess the effect of window size on the anomaly detection performance of our TSAD methods, we set it to 1, 3, 5, 10, 20, 50, 100, 200, 500 and 1000 to represent a wide range of possible values. The resulting figure is plotted in Figure 1.
@@
In this figure, we observe the F1 scores for various window sizes when running GDN on the SWAT dataset. The variability in the model’s performance across different window sizes is immediately noticeable. For instance, changing the window size from 1 to 3 results in more than a $5\%$ decrease in F1 score. This should not occur; in fact, we should expect the model to perform better when provided with more data within the window. It is also worth noting that even a $3\%$ improvement is considered groundbreaking in the literature. The sharp drop in performance for a window size of $500$ is also alarming. Although several trials suggest it may be an outlier, this is still far from ideal, as an anomaly detection model in practice should not have drastic dips in performance merely due to changing a parameter.
@@
Not only is there a lack of correlation between window size and performance, but using a window size of 1 actually yielded the highest F1 score in this experiment. This is unexpected because GDN is a predictive method, and a window size of 1 means the model only analyzes the previous time point when determining if a given time point is anomalous. This suggests that GDN on SWAT is not effectively capturing long-term dependencies when using larger window sizes; if it were, we would expect better performance for window sizes above 1, at least up to a certain point. We performed similar window size experiments on the MTAD-GAT and USAD models and observed the same trend. Our main takeaway is that these models do not sufficiently utilize their window size architecture to capture long-term dependencies in datasets such as SWAT. Additionally, they suffer from large performance variance relative to the window size, often much more than 5%. This variation in performance is far greater than what current methods achieve over traditional methods, challenging the reliability of their claims, as the observed performance boost might be attributed to changing a parameter such as window size. It is also worth noting that there is no consistency in the literature regarding the standard window size for a dataset. Furthermore, some papers claim to downsample the data due to memory concerns, which also impacts the amount of information in each window and can drastically affect performance compared to the existing SOTA methods.
III-B Seed Variability
@@
Many model initialization and data preprocessing steps involve randomness. To ensure the reproducibility of results, it is crucial to evaluate the impact of seed variability. Running experiments with different seeds helps in understanding the stability of the initialization methods and the algorithms’ performance. Ideally, changing the seed number should not significantly affect performance. If it does, it suggests that the method is not reliable and that optimal seed selection might have been done using grid search to inflate the model’s performance compared to established baselines. In our experiments, we ran the models with their default parameters but fixed the seed to various seed numbers, namely $s\in\{1,2,\ldots,25\}$, although any $25$ different seeds would work. The goal was to examine the variability of the models.
@@
Figure 2: F1 scores obtained by GDN and MTAD-GAT on the SWAT dataset for varying random seed numbers.
@@
Figure 2 illustrates 25 trials of GDN on SWAT with the exact same initial parameters. Among the first 10 seeds, there are sharp increases and decreases, many exceeding $5\%$ changes or more. This already supports our claim from the previous section that the model’s results are too variable and the models are hypersensitive to parameters, which should not be the case. Upon inspecting and plotting $25$ trials, we observe even more drastic changes in performance. In 3 of these 25 seeds, the model’s performance dropped below a 0.55 F1 score, representing more than a $25\%$ change from the other F1 values. This is very alarming, as we would expect a reliable model to be robust across multiple trials with the same input parameters. Figure 2 also displays the same trend but with the MTAD-GAT model: Some trials yield F1 scores above $0.8$ while others yield F1 scores below $0.3$. We observe once again that with no input parameter change, there is a discrepancy between the F1 and overall performance of MTAD-GAT across different trials. The results of these experiments reiterate the point raised in the previous section: current SOTA models seem to be overly sensitive to hyperparameters such as seed number or window size. This sensitivity can provide a shortcut to improved performance simply by searching for and reporting the best parameters for the model rather than for the baselines. We observe performance changes as high as double digits just by perturbing the window size or changing the seed number, which is far greater than the claimed improvements due to their innovative architecture compared to SOTA. This casts a shadow of doubt over which factor plays the major role in the observed improvements: the model itself or the initialization parameters.
III-C Normalization
@@
III-D Analysis of Anomaly Score
@@
As observed in previous sections, many seemingly unrelated parameters can significantly affect the performance of a TSAD model. As several studies have hinted earlier, the current numerical improvements many papers reported are often due to flawed thresholding and evaluation protocols. To have a more realistic view of the true output of TSAD models under different conditions, we analyze the density of raw anomaly scores before applying any transformation, thresholding, or point adjustment. We plot the density for training, validation, and test sets: in the SWAT and SMD datasets, the training and validation sets consist only of normal data, while the test set includes both normal and abnormal data. Ideally, we would like to observe significantly low anomaly scores for the training, validation, and normal test data, and significantly higher anomaly scores for the abnormal test data.
Figure 4 illustrates the anomaly score density of GDN on the SMD dataset. We can observe that the anomaly scores of the training samples are what we expect: near zero, since the model learned during training that these were normal. The validation set, which consists of unseen normal data, also shows anomaly scores close to zero. However, the density of the anomaly scores for abnormal samples during the test phase does not significantly differ from the normal data, and is still close to zero. This explains why the model does not perform well in this experimental setup: it cannot distinguish between normal and abnormal samples. Yet, by incorporating schemes such as point adjustment, it might be possible to report an acceptable F1 score. Another interesting and extreme example of counterintuitive anomaly score densities is observed with MTAD-GAT on the SWAT dataset. As demonstrated by the dashed lines in Figure 4, the anomaly scores of the normal test data are even larger than those of the abnormal data. This is counterintuitive given that the model is trained to minimize the anomaly score on normal data. However, since the density of normal test data scores does not fully overlap with the distribution of abnormal anomaly scores, the model could still achieve a good F1 score by using a kernel to transform the anomaly scores and thresholding to isolate the densities. This highlights how post-processing techniques can mask underlying model deficiencies and produce seemingly favorable results. At the cost of sacrificing the robustness of the model, one could thoroughly search for a set of optimal initialization parameters such as window size, seed number, and normalization technique. Combined with existing flawed evaluation protocols, this approach could potentially outperform any SOTA method without actually improving the model’s impact on real-world TSAD problems. This emphasizes the need for rigorous evaluation and transparent reporting of preprocessing steps and input data characteristics to ensure the reliability and fairness of anomaly detection methods.
@@
Figure 4: Density plots of anomaly scores obtained by GDN on the SMD dataset and MTAD on the SWAT dataset. The plot illustrates the distribution of anomaly scores across different sets: the training set, validation set, normal data from the test set, and abnormal data from the test set.
IV Conclusion